# Artificial intelligence for science: The easy and hard problems

### v1

@@ Abstract
Most work applying AI to science has focused on what might be called the “easy problem.” This is a relative term, since the easy problem is actually quite hard. A scientist specifies a function that they want to optimize (e.g., a function that generates a protein’s structure given its amino acid sequence). Included in the specification is the input for the function (e.g., the amino acid sequence), the output (e.g., the 3D structure), and a way to compare the function’s output with the ground truth (e.g., the average 3D distance of an amino acid residue from where it should be). The scientist then finds or collects a dataset, usually very large, with examples of the ground truth; or, designs some other way of assessing the model’s output (e.g., turbulence parameters in plasma flow). AI optimization tools can then be applied to this problem. So far, this kind of application has been highly successful, with new discoveries of tertiary protein structures, antibiotics, and nuclear fusion reactor designs (see [1] for a recent review).
@@
Are these conceptual breakthroughs just patterns that can be discovered with a sufficiently powerful pattern recognition system? In a sense yes, but before that can happen, something has to tell the pattern recognition system what kind of patterns are interesting, important, and useful. What problem is the pattern-recognition system designed to solve, and where does this come from?
@@
Much of the classic work on AI science (mainly by Simon, Langley, and their collaborators [3, 4, 5], but also more recently by Schmidt & Lipson [6], Udrescu & Tegmark [7], and others [8]) focused on the easy problem. For Simon and Langley, this approach was premised on the psychological thesis that scientific cognition was essentially the same as regular problem solving, only applied to a different (and sometimes more challenging) set of problems. Consequently, they developed algorithms that emulated human problem solving, and applied these to scientific discovery.
Existing AI scientists have had some success at the easy problem. Simon, Langley, and others were able to solve a range of seminal science problems [9], including the (re-)discovery of oxygen with STAHLp [10]; more modern methods for automated physics have inferred many existing and novel laws, including classical and quantum problems with AI Feynman and non-linear dynamical systems with SINDy [7, 8]; and, discovery algorithms in biology have advanced our ability to solve many difficult problems, including AlphaFold2 for protein folding [11].
This success is analogous to the earliest use of computers, in which they were used to complete calculations too laborious for any human (such as in the Enigma cryptography project in World War II; or, to prove all edge cases of a complicated theorem [12]). Algorithms that solve the easy problems of science are useful, even essential, to progress. For example, there is an increasing discrepancy between the number of amino-acid sequences that are discovered in biology and the recovery of the 3D protein structures they correspond to using the experimental method.
While recognizing the importance of such algorithms, we should also recognize their limitations. Several decades ago, Chalmers, French, & Hofstadter (focusing on the models of Simon, Langley, and their collaborators) challenged the idea that this kind of optimization was a complete model of scientific discovery and investigation [13]. Systems like STAHLp are only able to solve scientific problems and make discoveries, they argued, because the modelers have represented the inputs and outputs to the problem in hindsight; only relevant data have been included and those data are already organized such that the proposed heuristics will be able to easily extract the right solution. In other words, they have been provided a representation of the scientific problem that already includes the basic primitives needed for the final theory, but skirt the central problem of representation itself: Where do the primitives come from, and how do we know if we have discovered the right ones?
@@
In machine learning terms, these systems might be extremely good at interpolation, and they may become better at extrapolation to new data, but they will never automatically generate or choose to investigate new scientific problems. This is because neither the inferential nor the learned components of the algorithm contain the knowledge necessary to do so. Instead, that knowledge came from the team that specified the problem by choosing how to represent the inputs, outputs, and objective function.
Problem representation
The representation of input data involves two fundamental choices—which are the right primitive variables and which datapoints to include. In trying to emulate the investigative processes of scientists, it is important to consider the primitives they would have begun with. The representation chosen for the inputs cannot be too permissive—it cannot use concepts or data that scientists came up with in the course of solving the problem; nor too restrictive—it cannot exclude concepts or data that the would have originally affected the problem-solving process.
The output for a scientific problem comprises the scientific theory and any predictions it generates. This choice of representation determines which theories are considered “well-formed” and therefore valid solutions for the problem. The output representation can be defined either explicitly, in the form of a set of symbols and operations, or implicitly, through the space of operations that can be applied to the input variables. Modeling the full scientific process requires specifying a generative system for theories that has sufficient flexibility for conceptual change, which in turn might affect the space of theories that are considered well-formed.
The third component of a problem representation is the goal, expressed in the language of optimization as a loss function that assesses the adequacy of a solution compared to the “ground-truth.” The choice of loss function corresponds closely to the way the modeler has chosen to represent the structure of the natural domain. For example, when providing deterministic physical models for cosmically short distances, loss functions based on Euclidean distance might be suitable; for classification models, some assessment of decision accuracy; for probabilistic models, a loss based on relative entropy. The choice of loss is also influenced by the cognitive biases of scientists themselves. A classic finding from cognitive science is that for perceptual stimuli with separable feature dimensions participants’ generalizations are better captured with a Manhattan loss, in contrast to a Euclidean loss for stimuli with integral dimensions [15].
@@
In contemplating how to build AI systems that solve the hard problem, it is instructive to look at how human scientists do it. The high-level objective in science is clear: We would like to account for more data with our theories. At this level, human scientists break the hard problem down into several sub-problems:
@@
Once the domain and constraints have been specified, we can define an optimization problem (theory search); hence, we have converted the hard problem into the easy problem. For most current AI scientists, the modeling team conducts domain specification in advance in the representation and selection of data, and constraint specification in the representational scheme for potential scientific theories (outputs) and the objective function that assesses them. However, it is uncommon for real scientists to do a single pass from hard to easy, because they often realize that the problem they are solving is the wrong one. This may happen for several reasons. One is the realization that a theory is internally inconsistent or paradoxical. Another is the realization that the theory may (with suitable modification) be able to explain a broader range of phenomena, prompting a respecification of the domain. Conversely, phenomena which were previously included in a domain may need to be excluded if no adequate unifying theory is found for all the phenomena. Respecification can also happen when new empirical phenomena are reported. In a related vein, constraint respecification can happen when domains are merged, split, expanded, or shrunk. The key point is that problem creation and problem solving are cyclically coupled in scientific practice.
In the following sections, we motivate the distinction between the easy and hard problems with three case studies from the birth of modern chemistry, physics, and molecular biology. For each case study, we summarize the elements of the problem, the historical setting, and modern computational systems that have tried to recapture some aspects of these discoveries. We will argue that none of these modern systems offers a complete solution to the hard problem.
@@
In the 18th century, it had been observed that lead increased in weight when it was slowly heated (which today we call “oxidation,” but at that time was called “calcination”). This was difficult to explain with contemporary chemical theories, because they posited that something left a metal when it was heated (a type of inflammable earth called “phlogiston”). In 1774, the English chemist Joseph Priestley collected and identified a particularly inflammable and respirable form of air following the thermal reduction of calx-of-mercury (mercury-oxide) [16]. The French chemist Antoine Lavoisier eventually called this air “oxygen,” and posited that it went into the metal during calcination instead, causing the weight change [17]. Lavoisier’s course of investigations were so successful that he has been credited as having started the Chemical Revolution and introduced the principled application of the conservation of mass into the quantitative sciences.
@@
Rose and Langley proposed a computational model called STAHLp to account for the discovery of the role of oxygen in calcination reactions [10] (see Box 1).
@@
The input to STAHLp is a set of interconnected beliefs about 1) which substances are present before and after a particular reaction or 2) the chemical composition of each substance. These inputs are encoded using two types of variable: The functions (or programs) REACTS and COMPOSED OF, which operate on an unbounded space of discrete chemical names.
STAHLp’s desired output is a coherent and consistent “theory”—a set of beliefs that entail the inputs and do not contradict each other. STAHLp uses a hard objective function to enforce this: The theory cannot contain any inconsistent equations containing “nil” (the empty set).
STAHLp solves this problem by applying a set of “production rules” to its beliefs at each step, generating further beliefs. If the system generates an inconsistent belief, STAHLp throws an error. At that point, a second set of “belief revision” heuristics is applied to try to identify the source of the inconsistency and correct it. After the lowest-cost correction is made, STAHLp applies its production rules to generate the updated theory entailed by the new starting beliefs. The algorithm keeps running until no remaining beliefs are or an inconsistency has been detected; this is what Simon meant by scientific problem-solving as search through a hypothesis space [3].
@@
Rose and Langley showed that for a particular pair of beliefs STAHLp “discovers” oxygen (see Figure 1). In particular, the first belief, inherited from Georg Ernst Stahl, states that mercury is composed of calx-of-mercury and phlogiston. The second, reflecting an empirical observation by Joseph Priestley, states that calx-of-mercury is composed of mercury and a colorless gas.
When STAHLp’s production rules are applied to these observations, they produce an inconsistent belief—mercury-calx can be decomposed into itself, phlogiston, and oxygen (a circularity present in the two starting observations), which is then reduced to a statement containing nil on the left-hand side.
The update-belief rules are triggered, generating a set of “effect hypotheses” that “balance” the inconsistent belief (belief 4 in Figure 1):
  • (EH1) CM [Ph O] $\rightarrow$ CM Ph O; left: missing Ph and O
  • (EH2) CM [Ph] $\rightarrow$ CM Ph (O); left: missing Ph: right: extra O
  • (EH3) CM [O] $\rightarrow$ CM (Ph) O; left: missing O; right: extra Ph
  • (EH4) CM $\rightarrow$ CM (Ph O); right: extra Ph and O
By back-tracing where the left and right sides come from, STAHLp can generate “cause-hypotheses” about how the initials beliefs should be updated. The cause-hypothesis that affect the least downstream statements is chosen—in this case the addition of “oxygen” to the left hand side of belief 1.
Figure 1: The discovery of oxygen by STAHLp. The input system of two beliefs about the composition of mercury (M) and calx-of-mercury (CM), reflecting the knowledge in the early 1770s. Ph refers to phlogiston, and O to the air observed by Priestley that Lavoisier eventually named oxygen. Reproduced from [10].
STAHLp: Analysis
It is hard to argue with the assumption that reactions and compositions of substances were the central concepts in chemistry—indeed, this was how Stahl himself defined its scope [18]. However, leading up to the chemical revolution, chemists had a different way of thinking about the internal structure of substances, in which observable substances arose from mixtures of the latent primitives of earth, water, and fire. A new name could not be added arbitrarily, and had to be placed within the existing ontological structure. Second, they would not have considered air—what we now call gas—to have chemical properties and enter into chemical combinations: The discrete name representation is too permissive, and “oxygen” is simply not a valid entry [19].
It also excludes relevant data. Most of the scientific work leading up to the discovery of oxygen was concerned with the sensory properties of substances,111appearance, taste, feel and ductility and smell. where they came from,222the sea, the animal kingdom, mining, etc. and their weight. For example, there was actually a great deal of inconclusive or even negative evidence that metals apart from lead increased weight on calcination, detracting from the general statement that an air entered into metals. Similar arguments can be put forward for using a single function to represent a number of different reactions.
There is also the question of data selection. The creators of STAHLp include only two facts from the many heterogenous and often inconsistent observations and beliefs in 18th century chemistry. If they took into account others—for example, that when nitrous acid was poured on mercury colored vapors and fumes were given off—the model’s conclusions may well have changed.
Finally, the objective function for STAHLp is based on the detection of nil statements. Once again, this is a retrospective assumption that relies on the application of the conservation of mass. By contrast, in the 18th century it was widely held that substances could dissipate away to nothing—from diamond [19], to phlogiston itself [20]. Lavoisier had to create the right conceptual framework needed to support the use of equations in his investigations before using them.
Historical perspective: Lavoisier and the discovery of oxygen
Instead of being bound by a fixed type structure, Lavoisier made a number of conceptual innovations that were closely informed by the ontological structure of chemical knowledge [21]. The first was that “air” (what we would now call gas) could be involved in chemical reactions at all. Robert Boyle and others had offered a physical interpretation of air and derived various laws. But, as surprising as it sounds today, in Continental Europe air was not thought to enter into chemical combinations—it was not a chemical type. By including air in the “definitions of chemistry,” Lavoisier respecified the domain to include gross changes in air volume, and in turn explained the gross weight changes in calcination by the chemical fixation of air.
@@
Next, Lavoisier broadened his scope to include all operations that fix or release air [22], with the aim of tracing the flow of air and water through different coupled reactions in order to infer the chemical composition of a more complex substance (like chalk). This richer set of data, including previous analyses by the Scottish chemist Joseph Black, led to the development of quantitative models based on equations. Prior to Lavoisier, chemists categorized and weighed solids and liquids before and after reactions, but did not routinely measure the air surrounding these materials. So, their “equations” seldom balanced, and the conservation of mass was used more as a post-hoc and abstract principle rather than a tool for quantitative purposes. Lavoisier developed the conceptual machinery to represent a reaction in terms of the total weight of materials at the start and end, and in doing so established the loss function to be optimized—the inference of a consistent and useful set of equations. In other words, he constructed the right representation of the problem. This placed emphasis on the use of a density constant to relate changes in air volume to changes in weight.
@@
Discrepancies in subsequent experiments led Lavoisier to the conclusion that there must be different subtypes of air with different densities. This led to the development of new equipment to measure those densities, and ultimately the finding that the air of the atmosphere was in fact a composite of these subtypes, rather than an elemental root. He then showed that the reduction of calx-of-mercury with charcoal produced a different air (carbon monoxide and carbon dioxide) than the reduction of calx-of-mercury without charcoal, eventually calling the latter air “oxygen.” Lavoisier explained the differences between these two reactions by positing an underlying, potentially infinite range of chemical primitives that could take the familiar three states of matter depending on how much of the “matter of fire” was coupled with them. This was the beginning of the main Chemical Revolution—actually more of an inversion, in that the things previously considered elemental (earth, water, air, fire) were now considered complex, whereas previously complex things like carbon were now considered elemental.
@@
By the middle of the 19th century, Michael Faraday had published a set of discoveries and observations related to electromagnetic induction: A current could be generated in a conducting wire in the presence of a strong permanent magnet by moving the magnet or the wire. Faraday recorded the intensity of magnetic force surrounding magnets of various shapes, strengths, and number, as well as electrical circuits, arguing that the most useful representation for these data was in terms of lines of magnetic force [23]. He had speculated on what might be the cause of these patterns, but had been largely unsuccessful [24]. The Scottish physicist James Clerk Maxwell derived a brilliant and creative theoretical solution to this problem that provides the foundation of modern physics—the mathematical representation of the electromagnetic field.
@@
The input for AI Feynman is a data table, comprising data samples (rows) of a dependent variable and several independent variables (columns) that the modeler has specified in advance for each problem. Variables take continuous values, correspond to measurements of the physical system, and are augmented with type information representing their fundamental physical units (meter, second, kilogram, kelvin, and volt.).
AI Feynman outputs predictions that match the dimensionality and type structure of the input, as well as a symbolic formula representing a theory of the observed system. The objective function uses a squared-error loss to assess predictions in the input space and a hard loss on whether its current solution is equivalent to the ground-truth expression.
AI Feynman cycles through a set of computational strategies premised on commonalities in the functional forms of solutions to known physical problems (Figure Historical perspective: Maxwell and the electromagnetic field concept). The inputs and outputs to physical problems tend to have units, which justifies algebraic manipulations based on their types (dimensional analysis). Solutions, or parts thereof, often contain polynomial expressions, justifying polynomial fitting; they tend to be compositional, justifying search over symbolic expressions; they tend to be smooth, justifying approximation by a neural networks; they tend to exhibit symmetry and separability, allowing a reduction of variables after transformation by the neural network components. If nothing else works, a fixed set of transformations are applied to the variables, including the transcendental functions.
@@
For example, the data in “mystery table 5” comprises samples from one dependent variable, $F$, and nine independent variables corresponding to the masses and 3D positions of two objects, and Newton’s constant $G$. The algorithm runs through its pre-determined steps: Algebraic manipulations yield a reduced set of dimensionless variables; the application of a neural network component identifies translational symmetry; a good factorization is found; then polynomials are fit to two subsets of transformed variables. The end result of this process is an equation that accounts for the data below some error threshold, $\epsilon$ (see Figure Historical perspective: Maxwell and the electromagnetic field concept).
AI Feynman: Analysis
@@
Although the choice of input variables for AI Feynman might seem logical, they in fact correspond to quite an advanced stage of problem solving—when scientists have already constructed an idealized model for the system at hand. [^3] This is the process that Richard Feynman went through in his lectures when giving the historical background of the problem statement. For example, Newton had to posit the idea of a gravitational constant, expressed implicitly in terms of proportionality; and he had to posit that these were the only influential factors when explaining gravity—that action-at-a-distance was the correct framework to use, rather than the transmission of forces through an underlying medium. Similarly, Maxwell invented dimensional analysis to help solve difficult physics problems. But he did not always choose this representation—for electromagnetism, for example, he chose to think about dynamical properties of the aether.
There is also the question of which datapoints are chosen. For mystery table 5, the data are not taken from systems far from the scientist or near large masses, where the behavior of light (its speed or deflection, respectively) needs to be taken into account. Recognizing and adjusting for these factors were essential parts of proving the theory and then taking it forward.
Then there is the representation of theories. The choice of symbolic expressions to represent “natural laws” after the domain has been specified is reasonably unproblematic—although Newton himself did not use explicit symbols like “G” or formulae for the relationship between the motions of cosmic bodies [25]. Symbolic expressions are constrained to be “well-formed,” which requires that the modeling team ensures that each operation only runs on valid inputs and only produces valid outputs, defined by the initial primitives and the fixed type structure of the operations that run on them. Again, this scheme lacks the flexibility to capture the kinds of conceptual change that would have been necessary to derive the mature form of the problem. For example, the space of symbolic expressions might include primitives and operations that were discovered in the process of formulating the problem—analogous to providing the symbol $i$ before deriving a general solution to the problem of polynomial root finding. AI Feynman does not include primitives for the differential calculus, but the closely related algorithm SINDy does [8]. In this context, taking derivatives with non-integer powers might be required to capture the diffusion patterns in some data [26], but would not be allowed by its predefined type structure.
The kinds of laws AI Feynman can derive are also limited by its processing steps. This is motivated by an analysis of common characteristics of physical laws—they contain variables with units, low-degree polynomial structure, compositionality, smoothness, symmetry, and separability [7]. But again, these constraints arose out of analysis of the existing laws of physics, and provide constraints that restrict the subsequent class of models in an inflexible manner.
Historical perspective: Maxwell and the electromagnetic field concept
Nancy Nersessian has given a thorough cognitive-historical analysis of Maxwell and the development of the electromagnetic field concept [27]. In order to make progress given the ill-defined and heterogenous state of electrical science, Maxwell restricted his scope to Faraday’s data on electromagnetic induction and lines of force. In 1855, he gave a rigorous and analyzable form to Faraday’s observations and theoretical postulations using a descriptive mathematical model based on continuum mechanics of stresses in an underlying medium [28]. From 1861-1864, he tackled the deeper problem of providing a dynamical model that would explain these data [29, 30]. He began with magnetic phenomena, and showed that the constraints provided by his descriptive analysis could be fit by a vortex model. From this model he could calculate the magnetic force at any point in the medium by carrying over the system of equations describing the mechanical force exerted and replacing mechanical variables with magnetic ones [31].
When he generalized this model to a medium composed of these vortices, however, he found the model unsatisfactory because of the friction caused by adjacent vortices. This brought to mind the idle wheels interposed between rotating machine gears, from which he introduced the idea of idle-wheel-particles to communicate between vortices. Idle-wheel particles provided a good way to model electrical current, so his next step was to include electromagnetic phenomena. But this required the relaxation of the model to allow the particles to translate in conductive medium, and to rotate without generating any friction. Using the new model, he could bring in a set of equations to represent electrical current as the flux density of these particles, driven by the circumferential velocity of the vortices [31]. Maxwell continued this process of domain relaxation and model building to include electrostatic phenomena and the polarization of light.
@@
A striking feature of Maxwell’s problem solving is how explicit he was about the scope of his theories and the utility of intermediate models. Selectively restricting the domain allowed him to identify which parameters or features of the intermediate model were essential, and an analysis of those features afforded selective expansion of the domain—a process Nersessian has called “generic abstraction” [27]. Like Lavoisier, Maxwell was guided in this process of abstraction by ontological knowledge about the structure of different physical and mathematical systems, which also helped him sequentially assemble and modify the mathematical expressions underlying the model. Perhaps these idealized models played a role in Lavoisier’s early investigations, albeit in a simpler form involving crude movements of air and changes of weight. This process is not captured by systems like AI Feynman, which are given the problem variables from the mature idealized model, and lack the flexibility to alter their own conceptual systems.

### v2

@@ Abstract
Most work applying AI to science has focused on what might be called the “easy problem.” This is a relative term, since the easy problem is actually quite hard. A scientist specifies a function that they want to optimize (e.g., a function that generates a protein’s structure given its amino acid sequence). Included in the specification is the input for the function (e.g., the amino acid sequence), the output (e.g., the 3D structure), and a way to compare the function’s output with the ground truth or a desired range of outputs (e.g., the average 3D distance of an amino acid residue from where it should be). The scientist then finds or collects a dataset, usually very large, with examples of the ground truth, and applies AI optimization tools to the problem. So far, this kind of application has been highly successful, with new discoveries of tertiary protein structures, antibiotics, and nuclear fusion reactor designs (see [1] for a review).
@@
Much of the classic work on AI science (mainly by Simon, Langley, and their collaborators [3]; but also more recently [4, 5, 6, 7]) focuses on the easy problem. For Simon and Langley, this approach was premised on the psychological thesis that scientific cognition was essentially the same as regular problem solving, only applied to a different (and sometimes more challenging) set of problems. Consequently, they developed algorithms that emulated human problem solving, and applied these to model seminal scientific discoveries, including the (re-)discovery of oxygen with STAHLp [8]. More modern methods for automated physics have also inferred many existing and novel laws, including classical and quantum problems with AI Feynman and non-linear dynamical systems with SINDy [4, 5, 6]; and, discovery algorithms in biology have advanced our ability to solve many difficult problems, including AlphaFold2 for protein folding [7].
This success is analogous to the earliest use of computers, in which they were used to complete calculations too laborious for any human. Algorithms that solve the easy problems of science are useful, even essential, to progress. For example, there is an increasing discrepancy between the number of amino-acid sequences that are discovered in biology and the recovery of the 3D protein structures they correspond to using the experimental method.
While recognizing the importance of such algorithms, we should also recognize their limitations. Several decades ago, Chalmers, French, & Hofstadter challenged the idea that this kind of optimization was a complete model of scientific discovery and investigation [9]. Systems like STAHLp are only able to solve scientific problems and make discoveries, they argued, because the modelers have represented the inputs and outputs to the problem in hindsight. Only relevant data have been included and those data are already organized such that the proposed heuristics will be able to easily extract the right solution. In other words, most AI scientists have been provided a representation of the scientific problem that already includes the basic primitives needed for the final theory, but skirt the central problem of representation itself: Where do the primitives come from, and how do we know if we have discovered the right ones?
@@
In contemplating how to build AI systems that solve the hard problem, it is instructive to look at how human scientists do it. At a high level, humans break the hard problem down into two sub-problems:
@@
Once the domain and constraints have been specified, we can define an optimization problem (theory search); hence, we have converted the hard problem into the easy problem. However, it is uncommon for real scientists to do a single pass from hard to easy, because they often realize that the problem they are solving is the wrong one. This may happen for several reasons. One is the realization that a theory is internally inconsistent or paradoxical. Another is the realization that the theory may (with suitable modification) be able to explain a broader range of phenomena, prompting a respecification of the domain. Conversely, phenomena which were previously included in a domain may need to be excluded if no adequate unifying theory is found for all the phenomena. Respecification can also happen when new empirical phenomena are reported. In a related vein, constraint respecification can happen when domains are merged, split, expanded, or shrunk.
The key point is that problem creation and problem solving are cyclically coupled in scientific practice. For most current AI scientists, on the other hand, the modeling team has already conducted domain and constraint specification once, in advance—in the representation and selection of data, and the in the representational scheme for potential scientific theories (outputs) and the objective function that assesses them, respectively.
In the following sections, we motivate the distinction between the easy and hard problems with three case studies from the birth of modern chemistry, physics, and molecular biology. For each case study, we summarize the elements of the problem, the historical setting, and modern computational systems that have tried to recapture some aspects of these discoveries. We will argue that none of these systems offers a complete solution to the hard problem.
@@
In the 18th century, it had been observed that lead increased in weight when it was slowly heated (which today we call “oxidation,” but at that time was called “calcination”). This was difficult to explain with contemporary chemical theories, because they posited that something left a metal when it was heated (a type of inflammable earth called “phlogiston”). In 1774, the English chemist Joseph Priestley collected and identified a particularly inflammable and respirable form of air following the thermal reduction of calx-of-mercury (mercury-oxide) [12]. The French chemist Antoine Lavoisier eventually called this air “oxygen,” and posited that it went into the metal during calcination, causing the weight change [13]. Lavoisier’s course of investigations were so successful that he has been credited as having started the Chemical Revolution and introduced the principled application of the conservation of mass into the quantitative sciences.
@@
Rose and Langley proposed a computational model called STAHLp to account for the discovery of the role of oxygen in calcination reactions [8] (see Box 1). The input to STAHLp is a set of interconnected beliefs about 1) which substances are present before and after a particular reaction or 2) the chemical composition of each substance. These inputs are encoded using two types of variable: The predicates (or programs) REACTS and COMPOSED OF, and chemical names which the functions operate over. STAHLp’s desired output is a consistent “theory”—a set of beliefs that entail the inputs and do not contradict each other.
STAHLp solves this problem by applying a set of “production rules” to its beliefs, generating further beliefs until no more rules can be applied. If the system generates an inconsistent belief, STAHLp throws an error. At that point, a second set of “belief revision” heuristics is applied to try to identify the source of the inconsistency and correct it. After the lowest-cost correction is made, STAHLp applies its production rules to generate the updated theory entailed by the new starting beliefs. This is what Simon meant by scientific problem-solving as search through a hypothesis space [11].
Rose and Langley showed that for a particular pair of beliefs STAHLp “discovers” oxygen (see Figure 1). When STAHLp’s production rules are applied to these observations, they produce an inconsistent belief, which in turn triggers its update-belief rules, generating a set of “effect hypotheses” that “balance” the inconsistent belief (belief 4 in Figure 1). By back-tracing where the left and right sides come from, STAHLp can generate “cause-hypotheses” about how the initials beliefs should be updated to correct the inconsistency. The cause-hypothesis that affect the least downstream statements is chosen—in this case the addition of “oxygen” to the left hand side of belief 1.
Lavoisier
The modeling choices behind STAHLp neatly illustrate the easy vs hard distinction. It is hard to argue with the assumption that reactions and compositions of substances were the central concepts in chemistry—indeed, this was how Stahl himself defined its scope [14]. However, leading up to the chemical revolution, chemists had a different way of thinking about the internal structure of substances, in which observable substances arose from mixtures of the latent primitives of earth, water, and sometimes fire. Notably, most continental European chemists would not have considered air—what we now call gas—to have chemical properties and enter into chemical combinations [15]. By respecifying the “definitions of chemistry” to include air, Lavoisier was able to include measurements about gross changes in air volume in his theories, which in turn explained the gross weight changes during combustion, calcination, and reduction by the chemical fixation or release of air. This kind of conceptual innovation, through the revision of hierarchical ontological constraints, has been implicated at the core of human conceptual development [16].
@@
Figure 1: The discovery of oxygen by STAHLp. The input is two beliefs, where the first belief, inherited from Georg Ernst Stahl, states that mercury (M) is composed of calx-of-mercury (CM) and phlogiston (Ph). The second, reflecting an empirical observation by Joseph Priestley, states that calx-of-mercury is composed of mercury and a colorless gas (O). Through forward chaining the system reaches a circularity present in the initial beliefs—that mercury-calx can be decomposed into itself, phlogiston, and oxygen. The belief revision rules are then triggered, leading to a set of effect hypotheses, then cause hypotheses, giving the revisions needed to overcome the circularity. Reproduced from [8].
@@
Next, Lavoisier broadened his scope to include all operations that fix or release air [17], with the aim of tracing the flow of air and water through different coupled reactions in order to infer the chemical composition of a more complex substance (like chalk). This richer set of data about weight and volume changes led to the development of quantitative models based on tables and rudimentary equations. Prior to Lavoisier, chemists categorized and weighed solids and liquids before and after reactions, but did not routinely measure the air surrounding these materials. So, their “equations” seldom balanced, and the conservation of mass was used more as a post-hoc and abstract principle rather than a tool for quantitative purposes. In addition, in the 18th century it was widely held that substances could dissipate away to nothing—from diamond, to phlogiston itself [15].
@@
Discrepancies in subsequent experiments led Lavoisier to the conclusion that there must be different subtypes of air with different densities. This led to the development of new equipment to measure those densities, and ultimately the finding that the air of the atmosphere was in fact a composite of these subtypes, rather than an elemental root. He then showed that the reduction of calx-of-mercury with charcoal produced a different air (carbon monoxide and carbon dioxide) than the reduction of calx-of-mercury without charcoal, eventually calling the latter air “oxygen.” Lavoisier explained the differences between these two reactions by positing an underlying, potentially infinite range of chemical primitives that could take the familiar three states of matter depending on how much of the “matter of fire” was coupled with them. These primitives were simply those that could be isolated by the tools of the chemistry at the time. This was the beginning of the main Chemical Revolution—actually more of an inversion, in that the things previously considered elemental (earth, water, air, fire) were now considered complex, whereas previously complex things like carbon were now considered elemental.
With this historical framing, the it is easier to see how STAHLp’s specification of the problem is essentially a modern one. The objective function for STAHLp is based on the detection of nil statements, whereas Lavoisier had to develop the conceptual machinery—the right representation of the problem—to represent a reaction in terms of the total weight of materials at the start and end, and in doing so established the loss function to be optimized—the inference of a consistent and useful set of equations. In other words, he constructed the right representation of the problem. This placed emphasis on the use of a density constant to relate changes in air volume to changes in weight. Furthermore, a new chemical name could not be added arbitrarily, and had to be placed within the existing ontological structure that initially did not include air. “Oxygen” is simply not a valid entry [15]. Finally there is also the question of data selection. The creators of STAHLp include only two facts from the many heterogenous and often inconsistent observations and beliefs in 18th century chemistry. If they took into account others—for example, that when nitrous acid was poured on mercury colored vapors and fumes were given off—the model’s conclusions may well have changed. And, there was actually a great deal of inconclusive or even negative evidence that metals apart from lead increased weight on calcination, detracting from the general statement that an air entered into metals.
@@
In the mid-19th century, Michael Faraday published a set of discoveries and observations related to electromagnetic induction. Faraday recorded the intensity of current induced in a copper wire by moving it around magnets of various shapes, strengths, and number, as well as other electrical circuits embedded in magnetic media, arguing that the most useful representation for these data was in terms of lines of magnetic force [18]. He had speculated on what might be the cause of these patterns, but had been largely unsuccessful [19]. The Scottish physicist James Clerk Maxwell derived a brilliant and creative theoretical solution to this problem that provides the foundation of modern physics—the mathematical representation of the electromagnetic field.
@@
AI Feynman takes a data table, comprising data samples (rows) of a dependent variable and several independent variables (columns), and outputs a symbolic formula representing a theory of the system as well a set of predictions in the input space. Variables take continuous values, correspond to measurements of the physical system, and are augmented with type information representing their fundamental physical units (meter, second, kilogram, kelvin, and volt). In order to infer the theory of how the independent variables determine the dependent variable, AI Feynman cycles through a set of pre-specified computational strategies premised on commonalities in the functional forms of solutions to known physical problems (Figure Newton and Maxwell). The algorithm stops when the squared-error loss between its predictions and the input is low enough, and then checks whether its current solution is equivalent to the ground-truth expression.
@@
For example, the data in “mystery table 5” comprises samples from one dependent variable, $F$, and nine independent variables corresponding to the masses and 3D positions of two objects, and Newton’s constant $G$. The algorithm runs through its pre-determined steps: Algebraic manipulations yield a reduced set of dimensionless variables; the application of a neural network component identifies translational symmetry; a good factorization is found; then polynomials are fit to two subsets of transformed variables. The end result of this process is an equation that accounts for the data below some error threshold, $\epsilon$ (see Figure Newton and Maxwell).
Newton and Maxwell
@@
Although the input variables chosen for each problem in AI Feynman might seem logical, they in fact correspond to quite an advanced stage of problem solving—when scientists have already constructed an idealized model for the system at hand. [^1] This is actually the process that Richard Feynman goes through in his lectures when giving the historical background of the problem statement. For example, Newton had to posit the idea of a gravitational constant, expressed implicitly in terms of proportionality; and he had to posit that these quantities were the only influential factors when explaining gravity—that action-at-a-distance was the correct framework to use, rather than the transmission of forces through an underlying medium.
@@
Nancy Nersessian has given a thorough cognitive-historical analysis of Maxwell and the development of the electromagnetic field concept [20]. A striking feature of Maxwell’s problem-solving is how explicit he was about the scope of his theories and the utility of intermediate models. Selectively restricting the domain allowed him to identify which parameters or features of the intermediate model were essential, and an analysis of those features afforded selective expansion of the domain—a process Nersessian has called “generic abstraction.”
Maxwell began by restricting his scope to Faraday’s data on electromagnetic induction and lines of force, in order to make progress given the ill-defined and heterogenous state of electrical science at the time. First, he constructed a descriptive mathematical model of Faraday’s observations and theoretical postulations, based on the continuum mechanics of stresses in an underlying medium [21]. He then used this new representation of the data to design an explanatory dynamical model, iterating between adding to the range of phenomena he was considering, and adapting his intermediate model to accommodate them and generate new predictions [22, 23]. He began with magnetic phenomena, and showed that the constraints provided by his descriptive analysis could be fit by a vortex model. From this model he could calculate the magnetic force at any point in the medium by carrying over the system of equations describing the mechanical force that would be exerted in the vortex model and replacing mechanical variables with magnetic ones [24]. When he generalized this model to a medium composed of these vortices, however, he found the model unsatisfactory because of the friction that would occur between adjacent vortices. This brought to mind the idle wheels interposed between rotating machine gears, from which he introduced the idea of idle-wheel particles to communicate between vortices. Idle-wheel particles provided a good way to model electrical current, so his next step was to include electromagnetic phenomena. But this required the relaxation of the model to allow the particles to translate in conductive medium, and to rotate without generating any friction. Using the new model, he could bring in a set of equations to represent electrical current as the flux density of these particles, driven by the circumferential velocity of the vortices [24]. Maxwell continued this process of domain relaxation and model building to include electrostatic phenomena and the polarization of light.